# Learning with Incomplete Context: Linear Contextual Bandits with Pretrained Imputation

### v2

@@ Appendix B A data-driven approach for choosing $D_{t}$
$\left|(\bar{\bm{\theta}}_{t}-\widehat{\bm{\theta}}_{t})^{\top}\widehat{\bm{\phi}}_{t,A_{t}}\right|+\left|(\widehat{\bm{\theta}}_{t}-\bm{\theta}^{\star})^{\top}\widehat{\bm{\phi}}_{t,A_{t}}\right|$ $\leq 2\sqrt{\gamma_{t}\widehat{\bm{\phi}}_{t,A_{t}}^{\top}\bm{\Sigma}_{t}^{-1}\widehat{\bm{\phi}}_{t,A_{t}}}.$ (E.5)
@@
$\xi_{1,t}:=\min\left\{\sqrt{\gamma_{t}\widehat{\bm{\phi}}_{t,A_{t}}^{\top}\bm{\Sigma}_{t}^{-1}\widehat{\bm{\phi}}_{t,A_{t}}},1\right\}.$ (E.6)
@@
$\xi_{2,t}=\max_{a\in\mathcal{A}}\left|\bm{\theta}^{\star\top}\left(\widehat{\bm{\phi}}_{t,a}-\mathbb{E}\left[\bm{\Phi}\left(\bm{Y}_{t},a\right)\mid\bm{S}_{1:t}\right]\right)\right|.$ (E.7)
@@
$\xi_{2,t}\leq\max_{a\in\mathcal{A}}\left|\bm{\theta}^{\star\top}\left(\widehat{\bm{\phi}}_{t,a}-\mathbb{E}[\bm{\Phi}(\bm{Y}_{t},a)\mid\bm{S}_{1:t}]\right)\right|$ $\leq\left\|\bm{\theta}^{\star}\right\|_{2}\max_{a\in\mathcal{A}}\left\|\widehat{\bm{\phi}}_{t,a}-\mathbb{E}[\bm{\Phi}(\bm{Y}_{t},a)\mid\bm{S}_{1:t}]\right\|_{2}$ (E.8)
@@
Taking Equations (E.5), (E.6) and (E.7) into Equation (E.3), we have
$\left|\mathrm{reg}_{t}\right|$ $=\min\left\{\left|\mathrm{reg}_{t}\right|,1\right\}\leq 2\mathbb{E}\left[\xi_{1,t}\mid\bm{S}_{1:t}\right]+2\xi_{2,t},$ (E.9)
where the first equality follows from the assumption that $R(t,a)\in[-1,1]$ for any $a\in\mathcal{A}$. Summing Equation (E.9) over $t\in[T]$ gives
$\sum_{t=1}^{T}\mathrm{reg}_{t}$ $\leq 2\sum_{t=1}^{T}\mathbb{E}\left[\xi_{1,t}\mid\bm{S}_{1:t}\right]+2\sum_{t=1}^{T}\xi_{2,t}$ (E.10)
@@
where inequality $(i)$ follows from Cauchy-Schwarz inequality and inequality $(ii)$ follows from Equation (E.18) in Lemma E.1 and Equation (E.8).
@@
$R_{t}$ $=\bm{\theta}^{\star\top}\widehat{\bm{\phi}}_{t,A_{t}}+\bm{\theta}^{\star\top}\left(\mathbb{E}[\bm{\Phi}(\bm{Y}_{t},A_{t})\mid\bm{S}_{1:t},A_{t}]-\widehat{\bm{\phi}}_{t,A_{t}}\right)-\bm{\theta}^{\star\top}\left(\mathbb{E}[\bm{\Phi}(\bm{Y}_{t},A_{t})\mid\bm{S}_{1:t},A_{t}]-\bm{\Phi}(\bm{Y}_{t},A_{t})\right)+\eta_{t}$ (E.11)
@@
$\widehat{\bm{\theta}}_{t}-\bm{\theta}^{\star}$ $=\bm{\Sigma}_{t}^{-1}\sum_{\tau=1}^{t}R_{\tau}\widehat{\bm{\phi}}_{\tau,A_{\tau}}-\bm{\theta}^{\star}$ (E.12)
@@
where the first equality follows from Equation (E.11) and the last equality follows from the definition of $\Sigma_{t}$ in Equation (3.3).
@@
$\bm{\Sigma}_{t}^{-1}\sum_{\tau=1}^{t}\widehat{\bm{\phi}}_{\tau,A_{\tau}}\left(\mathbb{E}\left[\bm{\Phi}(\bm{Y}_{\tau},A_{\tau})\mid\bm{S}_{1:\tau},A_{\tau}\right]-\widehat{\bm{\phi}}_{\tau,A_{\tau}}\right)^{\top}\bm{\theta}^{\star}.$ (E.13)
Following the same analysis as Equation (E.8), we arrive at
$\left|\left(\mathbb{E}[\bm{\Phi}(\bm{Y}_{\tau},A_{\tau})\mid\bm{S}_{1:\tau},A_{\tau}]-\widehat{\bm{\phi}}_{\tau,A_{\tau}}\right)^{\top}\bm{\theta}^{\star}\right|\leq\left\|\mathbb{E}[\bm{\Phi}(\bm{Y}_{\tau},A_{\tau})\mid\bm{S}_{1:\tau},A_{\tau}]-\widehat{\bm{\phi}}_{\tau,A_{\tau}}\right\|_{2}\leq\sqrt{dD_{t}}.$ (E.14)
To control Equation (E.13), we have
$\left|\left(\sum_{\tau=1}^{t}\widehat{\bm{\phi}}_{\tau,A_{\tau}}^{\top}(\mathbb{E}[\bm{\Phi}(\bm{Y}_{\tau},A_{\tau})\mid\bm{S}_{1:\tau},A_{\tau}]-\widehat{\bm{\phi}}_{\tau,A_{\tau}})^{\top}\bm{\theta}^{\star}\right)\bm{\Sigma}_{t}^{-1}\left(\sum_{\tau=1}^{t}\widehat{\bm{\phi}}_{\tau,A_{\tau}}(\mathbb{E}[\bm{\Phi}(\bm{Y}_{\tau},A_{\tau})\mid\bm{S}_{1:\tau},A_{\tau}]-\widehat{\bm{\phi}}_{\tau,A_{\tau}})^{\top}\bm{\theta}^{\star}\right)\right|$ (E.15)
@@
where inequality $(i)$ follows from the triangle inequality, inequality $(ii)$ follows from Cauchy-Schwarz inequality and inequality $(iii)$ follows from Equation (E.14). Using properties of the trace operator, we continue to bound the right-hand side of Equation (E.15) using
$d\left(\sum_{\tau=1}^{t}D_{\tau}\right)\left(\sum_{\tau=1}^{t}\widehat{\bm{\phi}}^{\top}_{\tau,A_{\tau}}\bm{\Sigma}_{t}^{-1}\widehat{\bm{\phi}}_{\tau,A_{\tau}}\right)$ $=d\left(\sum_{\tau=1}^{t}D_{\tau}\right)\cdot\mathop{\mathrm{tr}}\left(\bm{\Sigma}_{t}^{-1}\sum_{\tau=1}^{t}\widehat{\bm{\phi}}_{\tau,A_{\tau}}\widehat{\bm{\phi}}^{\top}_{\tau,A_{\tau}}\right)$ (E.16)
@@
where equality $(i)$ follows from the definition of $\bm{\Sigma}_{t}$ as given in Equation (3.3). Taking Equation (E.16) into Equation (E.15) yields that
$\left\|\sum_{\tau=1}^{t}\widehat{\bm{\phi}}_{\tau,A_{\tau}}^{\top}\left(\mathbb{E}\left[\bm{\Phi}(\bm{Y}_{\tau},A_{\tau})\mid\bm{S}_{1:t},A_{\tau}\right]-\widehat{\bm{\phi}}_{\tau,A_{\tau}}\right)^{\top}\bm{\theta}^{\star}\right\|_{\Sigma_{t}^{-1}}\leq d\sqrt{\sum_{\tau=1}^{t}D_{\tau}}$ (E.17)
Therefore, using standard self-normalization concentration inequalities (see Lemma E.2) with Equations (E.12) and (E.17), with probability at least $1-\delta_{t}$,
@@
where the last inequality follows from Equation (E.21). It suffices to set $\delta_{t}:=\delta(3/\pi^{2})/t^{2}$. Hence, by taking $\gamma_{t}$ as defined in Equation (4.7), with probability at least $1-\delta$, we have
@@
holds for all $t\in[T]$. It follows from Equation (E.10) that $\sum_{t=1}^{T}\mathrm{reg}_{t}$ is bounded by
@@
E.1 Technical Lemmas
Lemma E.1.
For any $t\in[T]$ and $\xi_{1,t}$ defined in Equation (E.6), under the same conditions as Theorem 4.1, we have
$\sum_{t=1}^{T}\xi_{1,t}^{2}$ $\leq 2\gamma_{T}d\log\left(1+\frac{TB^{2}}{d\lambda}\right)$ (E.18)
@@ E.1 Technical Lemmas
$\sum_{t=1}^{T}\xi_{1,t}^{2}$ $\leq\sum_{t=1}^{T}\gamma_{t}\min\left\{\widehat{\bm{\phi}}_{t,A_{t}}^{\top}\bm{\Sigma}_{t}^{-1}\widehat{\bm{\phi}}_{t,A_{t}},1\right\}$ (E.19)
To control Equation (E.19), we use the potential function bound. We include a brief proof here for completeness. By the definition of $\bm{\Sigma}_{t+1}$ in Equation (3.3), we have
$\det\bm{\Sigma}_{t+1}$ $=\det\left(\bm{\Sigma}_{t}+\widehat{\bm{\phi}}_{t,A_{t}}\widehat{\bm{\phi}}^{\top}_{t,A_{t}}\right)=\det(\bm{\Sigma}_{t})\det\left(\bm{I}+\bm{\Sigma}^{-1/2}_{t}\widehat{\bm{\phi}}_{t,A_{t}}\left(\bm{\Sigma}_{t}^{-1/2}\widehat{\bm{\phi}}_{t,A_{t}}\right)^{\top}\right)$ (E.20)
@@
following Equation (E.20). Rearranging terms and taking logarithm on both sides of the above display implies that
$\sum_{t=1}^{T}\log\left(1+\widehat{\bm{\phi}}_{t,A_{t}}^{\top}\bm{\Sigma}_{t}^{-1}\widehat{\bm{\phi}}_{t,A_{t}}\right)=\log\left(\frac{\det\bm{\Sigma}_{T}}{\det\bm{\Sigma}_{0}}\right)\leq 2\gamma_{T}d\log\left(1+\frac{TB^{2}}{d\lambda}\right)$ (E.21)
where the last inequality follows from Assumption 2.1 and the potential function bound in Lemma E.3. Hence, applying the above display to Equation (E.19)
@@
where inequality $(i)$ follows from $\log(1+y)\geq y/2$ for all $y\in[0,1]$. Taking the above display into Equation (E.19) yields the desired bound as in Equation (E.18). ∎
Lemma E.2.
@@
See Theorem 1 in Abbasi-Yadkori et al., (2011). ∎
Lemma E.3 (Potential Function Bound).
@@
See Lemma 6.11 in Agarwal et al., (2019). ∎
Appendix F Proof of Results in Section 4.3
F.1 Proof of Proposition 4.1
@@
$\mathbb{E}_{\mathcal{D}}\left[\left\|\bm{b}-\widehat{\bm{b}}\right\|_{2}^{2}\right]\lesssim\frac{md_{S}}{NT_{0}},$ (F.1)
@@
$\mu_{t}:=\mathbb{E}\left[W_{t}\mid\bm{S}_{1:t}\right]=\sum_{j=0}^{m}\bm{\beta}_{j}^{\top}\bm{S}_{t-j}$ (F.2)
@@
$\widehat{\mu}_{t}:=\sum_{j=0}^{m}\widehat{\bm{\beta}}^{\top}_{j}\bm{S}_{t-j}$ (F.3)
@@
$\sqrt{D_{t}}$ $=\sqrt{\frac{1}{2}\operatorname{KL}\left(\mathcal{N}(\mu_{t},1)\|\mathcal{N}(\widehat{\mu}_{t},1)\right)}=\frac{1}{2}\left|\mu_{t}-\widehat{\mu}_{t}\right|$ (F.4)
@@
where the last equality follows from the definition of $\mu_{t}$ and $\widehat{\mu}_{t}$ in Equations (F.2). Since $\bm{S}_{t-j}$ and $\widehat{\bm{\beta}}_{j}$ are independent, conditioned on $\widehat{\bm{\beta}}_{j}$, we have
@@
Combining the above display with Equation (F.4) yields that
@@
where the last equality follows from Equation (F.1).
@@
$\mathbb{E}\left[\mathcal{R}_{T}\right]$ $\leq\delta T+\mathbb{E}\mathrm{reg}_{T}^{(\texttt{imp})}+\mathrm{reg}_{T}^{(\texttt{lin})}$ (F.5)
@@
Taking the above display into Equation (F.5) yields the desired result. ∎
F.2 Proof of Proposition 4.2
@@
From the classical nonparametric statistics literature, there exists an estimator $\widehat{f}$ (such as the kernel estimator, see Chapter 1 of Tsybakov,, 2008) of $f$ that satisfies
$\mathbb{E}_{\mathcal{D}}\left[\left\|\widehat{f}-f\right\|_{L_{2}}^{2}\right]\lesssim N^{-\frac{2\beta}{2\beta+d_{S}}},$ (F.6)
@@
Taking expectation over the historical data and combining Equation (F.6) with the above display, we have

### v3

@@ Appendix D A data-driven approach for choosing $D_{t}$
$\left|(\bar{\bm{\theta}}_{t}-\widehat{\bm{\theta}}_{t})^{\top}\widehat{\bm{\phi}}_{t,A_{t}}\right|+\left|(\widehat{\bm{\theta}}_{t}-\bm{\theta}^{\star})^{\top}\widehat{\bm{\phi}}_{t,A_{t}}\right|$ $\leq 2\sqrt{\gamma_{t}\widehat{\bm{\phi}}_{t,A_{t}}^{\top}\bm{\Sigma}_{t}^{-1}\widehat{\bm{\phi}}_{t,A_{t}}}.$ (G.5)
@@
$\xi_{1,t}:=\min\left\{\sqrt{\gamma_{t}\widehat{\bm{\phi}}_{t,A_{t}}^{\top}\bm{\Sigma}_{t}^{-1}\widehat{\bm{\phi}}_{t,A_{t}}},1\right\}.$ (G.6)
@@
$\xi_{2,t}=\max_{a\in\mathcal{A}}\left|\bm{\theta}^{\star\top}\left(\widehat{\bm{\phi}}_{t,a}-\mathbb{E}\left[\bm{\Phi}\left(\bm{Y}_{t},a\right)\mid\bm{S}_{1:t}\right]\right)\right|.$ (G.7)
@@
$\xi_{2,t}\leq\max_{a\in\mathcal{A}}\left|\bm{\theta}^{\star\top}\left(\widehat{\bm{\phi}}_{t,a}-\mathbb{E}[\bm{\Phi}(\bm{Y}_{t},a)\mid\bm{S}_{1:t}]\right)\right|$ $\leq\left\|\bm{\theta}^{\star}\right\|_{2}\max_{a\in\mathcal{A}}\left\|\widehat{\bm{\phi}}_{t,a}-\mathbb{E}[\bm{\Phi}(\bm{Y}_{t},a)\mid\bm{S}_{1:t}]\right\|_{2}$ (G.8)
@@
Taking Equations (G.5), (G.6) and (G.7) into Equation (G.3), we have
$\left|\mathrm{reg}_{t}\right|$ $=\min\left\{\left|\mathrm{reg}_{t}\right|,1\right\}\leq 2\mathbb{E}\left[\xi_{1,t}\mid\bm{S}_{1:t}\right]+2\xi_{2,t},$ (G.9)
where the first equality follows from the assumption that $R(t,a)\in[-1,1]$ for any $a\in\mathcal{A}$. Summing Equation (G.9) over $t\in[T]$ gives
$\sum_{t=1}^{T}\mathrm{reg}_{t}$ $\leq 2\sum_{t=1}^{T}\mathbb{E}\left[\xi_{1,t}\mid\bm{S}_{1:t}\right]+2\sum_{t=1}^{T}\xi_{2,t}$ (G.10)
@@
where inequality $(i)$ follows from Cauchy-Schwarz inequality and inequality $(ii)$ follows from Equation (G.18) in Lemma G.1 and Equation (G.8).
@@
$R_{t}$ $=\bm{\theta}^{\star\top}\widehat{\bm{\phi}}_{t,A_{t}}+\bm{\theta}^{\star\top}\left(\mathbb{E}[\bm{\Phi}(\bm{Y}_{t},A_{t})\mid\bm{S}_{1:t},A_{t}]-\widehat{\bm{\phi}}_{t,A_{t}}\right)-\bm{\theta}^{\star\top}\left(\mathbb{E}[\bm{\Phi}(\bm{Y}_{t},A_{t})\mid\bm{S}_{1:t},A_{t}]-\bm{\Phi}(\bm{Y}_{t},A_{t})\right)+\eta_{t}$ (G.11)
@@
$\widehat{\bm{\theta}}_{t}-\bm{\theta}^{\star}$ $=\bm{\Sigma}_{t}^{-1}\sum_{\tau=1}^{t}R_{\tau}\widehat{\bm{\phi}}_{\tau,A_{\tau}}-\bm{\theta}^{\star}$ (G.12)
@@
where the first equality follows from Equation (G.11) and the last equality follows from the definition of $\Sigma_{t}$ in Equation (3.3).
@@
$\bm{\Sigma}_{t}^{-1}\sum_{\tau=1}^{t}\widehat{\bm{\phi}}_{\tau,A_{\tau}}\left(\mathbb{E}\left[\bm{\Phi}(\bm{Y}_{\tau},A_{\tau})\mid\bm{S}_{1:\tau},A_{\tau}\right]-\widehat{\bm{\phi}}_{\tau,A_{\tau}}\right)^{\top}\bm{\theta}^{\star}.$ (G.13)
Following the same analysis as Equation (G.8), we arrive at
$\left|\left(\mathbb{E}[\bm{\Phi}(\bm{Y}_{\tau},A_{\tau})\mid\bm{S}_{1:\tau},A_{\tau}]-\widehat{\bm{\phi}}_{\tau,A_{\tau}}\right)^{\top}\bm{\theta}^{\star}\right|\leq\left\|\mathbb{E}[\bm{\Phi}(\bm{Y}_{\tau},A_{\tau})\mid\bm{S}_{1:\tau},A_{\tau}]-\widehat{\bm{\phi}}_{\tau,A_{\tau}}\right\|_{2}\leq\sqrt{dD_{t}}.$ (G.14)
To control Equation (G.13), we have
$\left|\left(\sum_{\tau=1}^{t}\widehat{\bm{\phi}}_{\tau,A_{\tau}}^{\top}(\mathbb{E}[\bm{\Phi}(\bm{Y}_{\tau},A_{\tau})\mid\bm{S}_{1:\tau},A_{\tau}]-\widehat{\bm{\phi}}_{\tau,A_{\tau}})^{\top}\bm{\theta}^{\star}\right)\bm{\Sigma}_{t}^{-1}\left(\sum_{\tau=1}^{t}\widehat{\bm{\phi}}_{\tau,A_{\tau}}(\mathbb{E}[\bm{\Phi}(\bm{Y}_{\tau},A_{\tau})\mid\bm{S}_{1:\tau},A_{\tau}]-\widehat{\bm{\phi}}_{\tau,A_{\tau}})^{\top}\bm{\theta}^{\star}\right)\right|$ (G.15)
@@
where inequality $(i)$ follows from the triangle inequality, inequality $(ii)$ follows from Cauchy-Schwarz inequality and inequality $(iii)$ follows from Equation (G.14). Using properties of the trace operator, we continue to bound the right-hand side of Equation (G.15) using
$d\left(\sum_{\tau=1}^{t}D_{\tau}\right)\left(\sum_{\tau=1}^{t}\widehat{\bm{\phi}}^{\top}_{\tau,A_{\tau}}\bm{\Sigma}_{t}^{-1}\widehat{\bm{\phi}}_{\tau,A_{\tau}}\right)$ $=d\left(\sum_{\tau=1}^{t}D_{\tau}\right)\cdot\mathop{\mathrm{tr}}\left(\bm{\Sigma}_{t}^{-1}\sum_{\tau=1}^{t}\widehat{\bm{\phi}}_{\tau,A_{\tau}}\widehat{\bm{\phi}}^{\top}_{\tau,A_{\tau}}\right)$ (G.16)
@@
where equality $(i)$ follows from the definition of $\bm{\Sigma}_{t}$ as given in Equation (3.3). Taking Equation (G.16) into Equation (G.15) yields that
$\left\|\sum_{\tau=1}^{t}\widehat{\bm{\phi}}_{\tau,A_{\tau}}^{\top}\left(\mathbb{E}\left[\bm{\Phi}(\bm{Y}_{\tau},A_{\tau})\mid\bm{S}_{1:t},A_{\tau}\right]-\widehat{\bm{\phi}}_{\tau,A_{\tau}}\right)^{\top}\bm{\theta}^{\star}\right\|_{\Sigma_{t}^{-1}}\leq d\sqrt{\sum_{\tau=1}^{t}D_{\tau}}$ (G.17)
Therefore, using standard self-normalization concentration inequalities (see Lemma G.2) with Equations (G.12) and (G.17), with probability at least $1-\delta_{t}$,
@@
where the last inequality follows from Equation (G.21). It suffices to set $\delta_{t}:=\delta(3/\pi^{2})/t^{2}$. Hence, by taking $\gamma_{t}$ as defined in Equation (4.7), with probability at least $1-\delta$, we have
@@
holds for all $t\in[T]$. It follows from Equation (G.10) that $\sum_{t=1}^{T}\mathrm{reg}_{t}$ is bounded by
@@
G.1 Technical Lemmas
Lemma G.1.
For any $t\in[T]$ and $\xi_{1,t}$ defined in Equation (G.6), under the same conditions as Theorem 4.1, we have
$\sum_{t=1}^{T}\xi_{1,t}^{2}$ $\leq 2\gamma_{T}d\log\left(1+\frac{TB^{2}}{d\lambda}\right)$ (G.18)
@@ G.1 Technical Lemmas
$\sum_{t=1}^{T}\xi_{1,t}^{2}$ $\leq\sum_{t=1}^{T}\gamma_{t}\min\left\{\widehat{\bm{\phi}}_{t,A_{t}}^{\top}\bm{\Sigma}_{t}^{-1}\widehat{\bm{\phi}}_{t,A_{t}},1\right\}$ (G.19)
To control Equation (G.19), we use the potential function bound. We include a brief proof here for completeness. By the definition of $\bm{\Sigma}_{t+1}$ in Equation (3.3), we have
$\det\bm{\Sigma}_{t+1}$ $=\det\left(\bm{\Sigma}_{t}+\widehat{\bm{\phi}}_{t,A_{t}}\widehat{\bm{\phi}}^{\top}_{t,A_{t}}\right)=\det(\bm{\Sigma}_{t})\det\left(\bm{I}+\bm{\Sigma}^{-1/2}_{t}\widehat{\bm{\phi}}_{t,A_{t}}\left(\bm{\Sigma}_{t}^{-1/2}\widehat{\bm{\phi}}_{t,A_{t}}\right)^{\top}\right)$ (G.20)
@@
following Equation (G.20). Rearranging terms and taking logarithm on both sides of the above display implies that
$\sum_{t=1}^{T}\log\left(1+\widehat{\bm{\phi}}_{t,A_{t}}^{\top}\bm{\Sigma}_{t}^{-1}\widehat{\bm{\phi}}_{t,A_{t}}\right)=\log\left(\frac{\det\bm{\Sigma}_{T}}{\det\bm{\Sigma}_{0}}\right)\leq 2\gamma_{T}d\log\left(1+\frac{TB^{2}}{d\lambda}\right)$ (G.21)
where the last inequality follows from Assumption 2.1 and the potential function bound in Lemma G.3. Hence, applying the above display to Equation (G.19)
@@
where inequality $(i)$ follows from $\log(1+y)\geq y/2$ for all $y\in[0,1]$. Taking the above display into Equation (G.19) yields the desired bound as in Equation (G.18). ∎
Lemma G.2.
@@
See Theorem 1 in Abbasi-Yadkori et al. (2011). ∎
Lemma G.3 (Potential Function Bound).
@@
See Lemma 6.11 in Agarwal et al. (2019). ∎
Appendix H Proof of Results in Section 4.3
H.1 Proof of Proposition 4.1
@@
$\mathbb{E}_{\mathcal{D}}\left[\left\|\bm{b}-\widehat{\bm{b}}\right\|_{2}^{2}\right]\lesssim\frac{md_{S}}{NT_{0}},$ (H.1)
@@
$\mu_{t}:=\mathbb{E}\left[W_{t}\mid\bm{S}_{1:t}\right]=\sum_{j=0}^{m}\bm{\beta}_{j}^{\top}\bm{S}_{t-j}$ (H.2)
@@
$\widehat{\mu}_{t}:=\sum_{j=0}^{m}\widehat{\bm{\beta}}^{\top}_{j}\bm{S}_{t-j}$ (H.3)
@@
$\sqrt{D_{t}}$ $=\sqrt{\frac{1}{2}\operatorname{KL}\left(\mathcal{N}(\mu_{t},1)\|\mathcal{N}(\widehat{\mu}_{t},1)\right)}=\frac{1}{2}\left|\mu_{t}-\widehat{\mu}_{t}\right|$ (H.4)
@@
where the last equality follows from the definition of $\mu_{t}$ and $\widehat{\mu}_{t}$ in Equations (H.2). Since $\bm{S}_{t-j}$ and $\widehat{\bm{\beta}}_{j}$ are independent, conditioned on $\widehat{\bm{\beta}}_{j}$, we have
@@
Combining the above display with Equation (H.4) yields that
@@
where the last equality follows from Equation (H.1).
@@
$\mathbb{E}\left[\mathcal{R}_{T}\right]$ $\leq\delta T+\mathbb{E}\mathrm{reg}_{T}^{(\texttt{imp})}+\mathrm{reg}_{T}^{(\texttt{lin})}$ (H.5)
@@
Taking the above display into Equation (H.5) yields the desired result. ∎
H.2 Proof of Proposition 4.2
@@
From the classical nonparametric statistics literature, there exists an estimator $\widehat{f}$ (such as the kernel estimator, see Chapter 1 of Tsybakov, 2008) of $f$ that satisfies
$\mathbb{E}_{\mathcal{D}}\left[\left\|\widehat{f}-f\right\|_{L_{2}}^{2}\right]\lesssim N^{-\frac{2\beta}{2\beta+d_{S}}},$ (H.6)
@@
Taking expectation over the historical data and combining Equation (H.6) with the above display, we have